# Formal Foundations for Controlled Stochastic Activity Networks

### v1

@@ Abstract
We develop a hierarchical, automata-theoretic semantics for Controlled SANs that encompasses nondeterministic, probabilistic, and stochastic models in a uniform manner. A structured taxonomy of control policies—ranging from memoryless and finite-memory strategies to computationally augmented policies—is formalized, and their expressive power is characterized through associated language classes. To support model abstraction and compositional reasoning, we introduce behavioral equivalences including bisimulation and stochastic isomorphism.
@@ 1 Introduction
Safety-critical systems—such as those used in avionics, autonomous vehicles, medical devices, and industrial automation—are characterized by stringent correctness requirements, continuous operation, and the need to make timely decisions under uncertainty. These systems typically involve a distributed architecture with interacting components that operate concurrently and must respond predictably to a combination of real-time stimuli, probabilistic events, and external commands. Any error in their behavior can have catastrophic consequences, making formal modeling and verification indispensable to their development and deployment [hoare1969axiomatic, lamport1977proving, newcombe2015aws, woodcock2009survey, seshia2022cyberphysical].
To analyze and synthesize such systems, we require modeling frameworks that are expressive enough to capture their complex dynamic behavior, yet structured enough to support rigorous reasoning and tool-based analysis. Classical paradigms—such as finite-state machines [Sipser2013], timed automata [alur1994theory], Petri nets [Peterson1981], and stochastic process algebras [hillston1996compositional, hermanns2002interactive]—have each contributed key insights in this direction. In particular, Stochastic Activity Networks (SANs) [movaghar1984stochastic, meyer1985stochastic, movaghar2001stochastic] generalize timed Petri nets by associating activities with probabilistic durations and gate-driven transitions. SANs are equipped with intuitive graphical notations and support compositionality, making them effective for performance and dependability analysis.
@@
We develop Controlled SANs through a layered approach that introduces increasing levels of semantic detail. At each level, we formalize system behavior using corresponding automata-theoretic models, define classes of control policies, and characterize the expressive power of the framework in terms of language recognition and bisimulation [Milner1989, baier2008principles].
@@
2. Probabilistic Semantics. To handle uncertainty in instantaneous transitions, we introduce controlled probabilistic activity networks, where the choice among enabled instantaneous activities in an unstable marking is made probabilistically. These are interpreted as controlled probabilistic automata, which generalize discrete-time Markov decision processes (DTMDPs) [puterman1994markov] with activity labels and control actions. This extension supports modeling fault injection, randomized algorithms, and probabilistic control behavior.
3. Stochastic Semantics. For real-time systems, we enrich the model with timing semantics by assigning probability distributions and enabling rates to timed activities. The resulting controlled stochastic activity networks are semantically equivalent to controlled stochastic automata, which model system dynamics as continuous-time stochastic processes governed by policies. In the special case where timing distributions are exponential, the models correspond to continuous-time Markov decision processes (CTMDPs) [puterman1994markov], enabling application of well-established analysis techniques.
@@
The rest of the paper is organized as follows. Section 2 discusses the related work. Section 3 introduces the syntax and semantics of Controlled SANs in a nondeterministic setting, along with their interpretation via controlled automata. Section 4 extends the model to controlled probabilistic activity networks, discussing policy types and language classes. Section 5 presents the stochastic semantics, introduces controlled stochastic automata, and formalizes the transition to CTMDPs. Section 6 summarizes the key findings and outlines future directions, including verification techniques and safety assurance methods for real-world applications.
@@ 2 Related Work
Stochastic Activity Networks (SANs) [movaghar1984stochastic, meyer1985stochastic, movaghar2001stochastic] extend timed Petri nets by associating activities with probabilistic timing and gate-based control logic. SANs have been extensively applied to performance modeling of distributed and fault-tolerant systems. However, they lack mechanisms for encoding control policies—that is, dynamic selection among multiple transitions based on system state or history. Our Controlled SANs generalize SANs by integrating control actions and policy-based semantics, subsuming the original SAN model as a special case without decision-making.
@@
The formal underpinnings of SANs are rooted in Petri net theory [Peterson1981], which has also been enriched through control-oriented extensions such as inhibitor arcs and priority transitions, and through the application of supervisory control theory to Petri nets [Ramadge1987, Cassandras2008]. These approaches often address control indirectly or externally. In contrast, Controlled SANs embed policy-driven behavior directly into the model’s semantics, enabling explicit and systematic reasoning about system evolution under control.
CTMDPs and GSMPs
Continuous-Time Markov Decision Processes (CTMDPs) [puterman1994markov] represent one of the canonical models for combining stochastic timing with nondeterministic control. They support optimal policy synthesis under quantitative objectives and are widely used in model checking and dependability analysis. Controlled SANs subsume CTMDPs in the special case where all timed activities follow exponential distributions and control actions determine probabilistic transitions. However, our framework allows richer structural and semantic features, including gate-driven transitions, compositionality, and generalized timing semantics.
Generalized Semi-Markov Processes (GSMPs) [Matthes1962, Schassberger1978] extend CTMDPs to support multiple concurrently enabled transitions with arbitrary distributions. While expressive, GSMPs lack an integrated control structure. Controlled SANs may be viewed as a control-augmented subclass of GSMPs, with a more explicit policy-theoretic semantics and better alignment with formal methods.
Timed Automata and Stochastic Hybrid Systems
Timed automata [alur1994theory] and their probabilistic and stochastic extensions [baier2008principles] offer a well-established formal framework for the specification and analysis of systems with real-time constraints. These models have been central to the development of verification tools such as UPPAAL and PRISM. However, traditional timed automata are clock-driven and best suited for deterministic or nondeterministic behaviors. Controlled SANs, in contrast, are activity-driven, using distributed timers with stochastic durations and gate-based enabling, which better captures concurrent probabilistic dynamics.
Stochastic Hybrid Systems (SHS)  [Cassandras2008, Tomlin1998] and Piecewise Deterministic Markov Processes (PDMPs)  [Davis1984] offer continuous-time dynamics with switching behaviors. While these models are more general in some respects, they often lack the compositional and language-theoretic semantics provided by automata-based models. Controlled SANs bridge this gap by supporting a hierarchy of semantics—from nondeterministic to stochastic—within a single automata-theoretic framework.
@@
The study of controlled automata and their language-theoretic properties under various policy types has a rich history in formal methods, including supervisory control theory [Ramadge1987], game-theoretic controller synthesis [Pnueli1989], and probabilistic automata [Segala1995]. Our work builds on these foundations by systematically classifying control policies (e.g., memoryless, finite-memory, stack-augmented, tape-augmented) and characterizing their expressive power in terms of accepted languages over finite and infinite words.
The hierarchy of language families induced by these policies mirrors classical distinctions in automata theory (e.g., regular, context-free, and recursively enumerable languages) and connects to foundational results on $\omega$-automata and Büchi acceptance [Thomas1990, Vardi1986Automata, Rabin1969].
@@
More general classes, such as stochastic BPA games [brazdil2010qualitative] and probabilistic pushdown automata (pPDA) [etessami2005recursive, etessami2009recursive, esparza2014probabilistic], capture recursion, nondeterminism, and probabilistic branching. These models correspond to infinite-state Markov decision processes with well-defined semantics and decidable reachability and $\omega$-regular properties for important subclasses.
@@
Finally, formal modeling and verification of safety-critical systems has increasingly emphasized the integration of stochastic timing, uncertainty, and control [seshia2022cyberphysical, Clarke1999]. Tools like PRISM [Kwiatkowska2011], MODEST [Hahn2010], and Storm [Dehnert2017] support a subset of these features, often limited to MDP-based models. Controlled SANs provide a formalism that aligns more closely with the practical demands of such systems—offering compositionality, distributed semantics, and structured control—while retaining formal rigor and support for verification and synthesis. One of the most prominent tools for modeling and analyzing SAN-based systems is the Möbius Modeling Framework [Deavours2002Mobius], which provides a compositional, extensible environment for performance and dependability analysis. Möbius has significantly influenced the modeling community and offers a natural foundation for extensions such as Controlled Stochastic Activity Networks (CSANs).
@@ 3 Nondeterministic Models
Controlled activity networks are a generalization of activity networks [movaghar1984stochastic, movaghar2001stochastic], incorporating control mechanisms in a nondeterministic setting.
@@
Except for the inclusion of control actions, the structure of controlled activity networks is the same as the structure of activity networks [movaghar2001stochastic]. As an example, consider the graphical representation of a controlled activity network as in Figure 1. $P1$, $P2$, $P3$, $P4$, $P5$, $P6$, $P7$, and $P8$ are places. $T1$, $T2$, $T3$, and $T4$ are timed activities, and $I1$, and $I2$ are instantaneous activities. $c1$, $c2$, $c3$, and $c4$ are control actions. $G1$ is an input gate with only one input, and $G2$ is an output gate with only one output. The enabling predicates, if any, and functions of these gates are indicated in a table called a “Gate Table” as depicted in Figure 1.
@@
Next, we define a notion of equivalence for controlled automata based on the concept of bisimulation, which is similar to the one proposed for automata [Milner1989, baier2008principles].
@@
The proof relies on a notable result about Petri nets—specifically, that the class of extended Petri nets [Peterson1981], which includes inhibitor arcs, can simulate nondeterministic Turing machines [Hack1975, Sipser2013]. Consequently, an activity network that incorporates instantaneous activities, standard gates, and inhibitor gates can likewise simulate a nondeterministic Turing machine and, therefore, represent any computable relation. This result is instrumental in the proof of the theorem. For the full proof, see Appendix A. ∎
The above proof shows that, in terms of behavioral equivalence, any system modeled by a computable controlled activity network can also be represented using a model that includes only standard gates and inhibitor gates. However, achieving this representation may require a significantly larger and more complex model, potentially at the cost of modeling clarity and convenience. Formally, this leads to the following corollary:
@@
Any computable controlled activity network is isomorphic to a controlled activity network that has only standard and inhibitor gates.
@@
We now define three types of policies for controlled automata by augmenting these models with finite-state automata, pushdown automata, and Turing machines [Sipser2013] as follows.
@@
$L_{0}(S)\subset L_{F}(S)\subset L_{\text{Stack}}(S)\subset L_{\text{Tape}}(S)\subset L_{H}(S)$
Proof Sketch..
This result follows directly from the arguments in the proof sketch of Theorem 3.29. ∎
@@
Proposition 3.31.
@@
Definition 3.32 (Uniform and Standard Closure under Language Operations).
@@
The next proposition shows that, at the regular level ($\Pi\in\{0,F\}$), this stronger form of uniform closure indeed holds via the standard disjoint-union and synchronous-product constructions.
Proposition 3.38.
@@
Proof Sketch.
For $\Pi\in\{0,F\}$, controlled automata reduce to nondeterministic finite automata (NFAs). The emptiness problem for NFAs can be decided in polynomial time via a reachability check over the underlying state graph [Hopcroft2007].
For $\Pi=\textit{Stack}$, the model corresponds to a pushdown automaton, and emptiness is known to be decidable using standard PDA algorithms [Hopcroft2007].
For $\Pi\in\{\textit{Tape},H\}$, the policy mechanism can simulate a Turing machine, and the emptiness problem reduces to the halting problem, which is undecidable [Sipser2013, Hopcroft1979]. ∎
@@
$L^{\omega}_{0}(S_{\omega})\subset L^{\omega}_{F}(S_{\omega})\subset L^{\omega}_{\text{Stack}}(S_{\omega})\subset L^{\omega}_{\text{Tape}}(S_{\omega})\subset L^{\omega}_{H}(S_{\omega})$
@@
As in the finite case, there is a strict hierarchy of global language families:
@@
$\mathbb{L}^{\omega}_{0}\subset\mathbb{L}^{\omega}_{F}\subset\mathbb{L}^{\omega}_{\text{Stack}}\subset\mathbb{L}^{\omega}_{\text{Tape}}=\mathbb{L}^{\omega}_{H^{\text{comp}}}\subseteq\mathbb{L}^{\omega}_{H}\,.$
Proof Sketch..
The argument parallels that of Theorem 3.29 for finite-word languages, adapted to $\omega$-languages under Büchi acceptance. Each inclusion corresponds to a strictly increasing level of controller power: memoryless $\subset$ finite-memory $\subset$ stack-augmented $\subset$ tape-augmented == computable history-dependent $\subset$ arbitrary history-dependent. Strictness follows from analogous examples where finite memory distinguishes finite phases, stacks enforce well-nested obligations, and tapes simulate arbitrary computations. Hence,
$\mathbb{L}^{\omega}_{0}\subset\mathbb{L}^{\omega}_{F}\subset\mathbb{L}^{\omega}_{\text{Stack}}\subset\mathbb{L}^{\omega}_{\text{Tape}}=\mathbb{L}^{\omega}_{H^{\text{comp}}}\subseteq\mathbb{L}^{\omega}_{H}\,.$
@@
Proposition 3.53.
@@
Proof Sketch..
This argument is the $\omega$-word analogue of Proposition 3.31. For each integer $k\geq 1$, define the $\omega$-regular language $W_{k}=(a^{k}b)^{\omega}$, i.e., the ultimately periodic infinite word consisting of blocks of $a^{k}$ followed by $b$, repeated forever. For distinct $k$, these $W_{k}$ are distinct $\omega$-regular languages.
@@
These results parallel Lemma 3.30 and Proposition 3.31 for finite words. The class $\mathbb{L}^{\omega}_{F}$ is strictly between $\mathbb{L}^{\omega}_{0}$ and $\mathbb{L}^{\omega}_{\mathrm{Stack}}$ and inherits the expressive power and closure properties of $\omega$-regular languages.
@@
Definition 3.55 (Uniform and Standard Closure under $\omega$-Language Operations).
@@
Proof Sketch.
By Büchi’s theorem [Buchi1962] and McNaughton’s determinization result [McNaughton1966], every $\omega$-regular language is accepted by a deterministic Muller automaton, and deterministic Muller automata are effectively closed under union, intersection, and complementation. Hence, $\mathrm{REG}^{\omega}$ enjoys full Boolean closure. ∎
@@
Proof Sketch.
Closure under union follows from the characterization of $\omega$-CFLs as finite unions of languages of the form $UV^{\omega}$ with $U,V$ context-free, as shown by Cohen and Gold [CohenGold1977]. Non-closure under intersection and complement is classical: there exist $\omega$-CFLs $L_{1},L_{2}$ such that $L_{1}\cap L_{2}$ is not $\omega$-context-free, implying, by De Morgan’s laws, failure of closure under complement (see also Staiger [Staiger1997] and Finkel [Finkel1997]). ∎
@@
Proof Sketch.
The class $\mathrm{RE}^{\omega}$ coincides with the family of analytic ($\Sigma^{1}_{1}$) subsets of $A^{\omega}$, known to be closed under finite union and intersection but not under complement (Kechris [Kechris1995], Moschovakis [Moschovakis2009]). ∎
@@
If, in addition, $\mathcal{S}_{\omega}$ is restricted to controlled deterministic parity automata (or any equivalent complement-closed $\omega$-regular formalism), then the same framework yields uniform closure under complementation as well, by replacing the parity condition with its dual on a fixed deterministic plant. Since we do not impose this structural restriction in general, Proposition 3.60 is stated only for union and intersection. It is well known that deterministic Büchi automata are not closed under complement, whereas deterministic parity or Rabin automata are. Hence, uniform complementation requires augmenting the acceptance structure beyond the Büchi property (e.g., to parity, Rabin/Street, or Muller conditions [Safra1988, Rabin1969, FrancezPnueli1980, Muller1963]).

### v2

@@ Abstract
We develop a hierarchical, automata-theoretic semantics for Controlled SANs that encompasses nondeterministic, probabilistic, and stochastic models uniformly. A structured taxonomy of control policies—ranging from memoryless and finite-memory strategies to computationally augmented policies—is formalized, and their expressive power is characterized through associated language classes. To support model abstraction and compositional reasoning, we introduce behavioral equivalences, including bisimulation and stochastic isomorphism.
@@ 1 Introduction
Safety-critical systems—such as those used in avionics, autonomous vehicles, medical devices, and industrial automation—are characterized by stringent correctness requirements, continuous operation, and the need to make timely decisions under uncertainty. These systems typically involve a distributed architecture with interacting components that operate concurrently and must respond predictably to a combination of real-time stimuli, probabilistic events, and external commands. Any error in their behavior can have catastrophic consequences, making formal modeling and verification indispensable to their development and deployment [28, 36, 52, 75, 63].
To analyze and synthesize such systems, we require modeling frameworks that are expressive enough to capture their complex dynamic behavior, yet structured enough to support rigorous reasoning and tool-based analysis. Classical paradigms—such as finite-state machines [65], timed automata [1], Petri nets [54], and stochastic process algebras [27, 25]—have each contributed key insights in this direction. In particular, Stochastic Activity Networks (SANs) [48, 44, 49] generalize timed Petri nets by associating activities with probabilistic durations and gate-driven transitions. SANs are equipped with intuitive graphical notations and support compositionality, making them effective for performance and dependability analysis.
@@
We develop Controlled SANs through a layered approach that introduces increasing levels of semantic detail. At each level, we formalize system behavior using corresponding automata-theoretic models, define classes of control policies, and characterize the expressive power of the framework in terms of language recognition and bisimulation [45, 3].
@@
2. Probabilistic Semantics. To handle uncertainty in instantaneous transitions, we introduce controlled probabilistic activity networks, where the choice among enabled instantaneous activities in an unstable marking is made probabilistically. These are interpreted as controlled probabilistic automata, which generalize discrete-time Markov decision processes (DTMDPs) [56] with activity labels and control actions. This extension supports modeling fault injection, randomized algorithms, and probabilistic control behavior.
3. Stochastic Semantics. For real-time systems, we enrich the model with timing semantics by assigning probability distributions and enabling rates to timed activities. The resulting controlled stochastic activity networks are semantically equivalent to controlled stochastic automata, which model system dynamics as continuous-time stochastic processes governed by policies. In the special case where timing distributions are exponential, the models correspond to continuous-time Markov decision processes (CTMDPs) [56], enabling application of well-established analysis techniques.
@@
The rest of the paper is organized as follows. Section 2 discusses the related work. Section 3 introduces the syntax and semantics of Controlled SANs in a nondeterministic setting, along with their interpretation via controlled automata. Section 4 extends the model to controlled probabilistic activity networks and discusses policy types and language classes. Section 5 presents the stochastic semantics, introduces controlled stochastic automata, and formalizes the transition to CTMDPs. Section 6 summarizes the key findings and outlines future directions, including verification techniques and safety assurance methods for real-world applications.
@@ 2 Related Work
Stochastic Activity Networks (SANs) [48, 44, 49] extend timed Petri nets by associating activities with probabilistic timing and gate-based control logic. SANs have been extensively applied to performance modeling of distributed and fault-tolerant systems. However, they lack mechanisms for encoding control policies—that is, dynamic selection among multiple transitions based on system state or history. Our Controlled SANs generalize SANs by integrating control actions and policy-based semantics, subsuming the original SAN model as a special case without decision-making.
A closely related and influential extension of Petri nets is the model of Generalized Stochastic Petri Nets (GSPNs) [40, 4]. GSPNs enrich classical Petri nets with both immediate (zero-time) and timed (transitionally stochastic) firings, yielding an underlying continuous-time Markov chain semantics. They provide a powerful framework for performance and dependability evaluation and have been widely adopted in the modeling community.
While GSPNs support rich stochastic timing and priority mechanisms, they do not incorporate explicit decision-making or policy-based control. In particular, nondeterminism in GSPNs is typically resolved implicitly (e.g., via priorities or probabilistic weights), rather than through externally selectable control actions. Controlled SANs extend beyond GSPNs by embedding control policies directly into the semantic model, enabling systematic reasoning about optimal decision-making, policy synthesis, and expressiveness under control.
@@
The formal underpinnings of SANs are rooted in Petri net theory [54], which has also been enriched through control-oriented extensions such as inhibitor arcs and priority transitions, and through the application of supervisory control theory to Petri nets [59, 9]. These approaches often address control indirectly or externally. In contrast, Controlled SANs embed policy-driven behavior directly into the model’s semantics, enabling explicit and systematic reasoning about system evolution under control.
CTMDPs, GSMPs, and GSMDPs
Continuous-Time Markov Decision Processes (CTMDPs) [56] provide a foundational model for decision-making under stochastic timing with nondeterministic control. They assume exponential holding times and enable optimal policy synthesis with respect to expected reward, reachability, or long-run behavior. Controlled SANs (CSANs) subsume CTMDPs when all timed activities are exponential, and control actions govern the resolution of nondeterminism, while extending the model with additional structure and semantic richness, including gate-based synchronization and compositional modeling.
Generalized Semi-Markov Processes (GSMPs) [42, 61] generalize CTMDPs by supporting multiple concurrently enabled transitions with arbitrary timing distributions and enabling race semantics. However, they are fundamentally descriptive and lack a native control-theoretic framework. CSANs may be seen as a control-augmented extension of GSMPs, with better alignment with formal verification and model-based synthesis.
Generalized Semi-Markov Decision Processes (GSMDPs) [76] introduce explicit decision-making into the GSMP framework, combining general timing distributions with controlled actions. These models are highly expressive and support value computation via phase-type approximation. However, GSMDPs often lack a compositional semantics, and their modeling language does not naturally capture structured constructs such as gates or multi-level synchronizations. In contrast, CSANs offer a modular, automata-based formalism that preserves analyzability while encompassing the timing and control features of both GSMPs and GSMDPs. In particular, the Controlled SAN model captures not only semi-Markovian timing and nondeterministic choice, but also allows interaction via gates, enabling a more expressive and implementable modeling paradigm.
From a semantic perspective, GSPNs can be viewed as a Petri-net–based counterpart of GSMPs, in which multiple concurrently enabled stochastic events race according to general timing distributions, but without explicit control actions.
@@
The study of controlled automata and their language-theoretic properties under various policy types has a rich history in formal methods, including supervisory control theory [59], game-theoretic controller synthesis [55], and probabilistic automata [62]. Our work builds on these foundations by systematically classifying control policies (e.g., memoryless, finite-memory, stack-augmented, tape-augmented) and characterizing their expressive power in terms of accepted languages over finite and infinite words.
The hierarchy of language families induced by these policies mirrors classical distinctions in automata theory (e.g., regular, context-free, and recursively enumerable languages) and connects to foundational results on $\omega$-automata and Büchi acceptance [69, 72, 58].
@@
More general classes, such as stochastic BPA games [7] and probabilistic pushdown automata (pPDA) [18, 19, 17], capture recursion, nondeterminism, and probabilistic branching. These models correspond to infinite-state Markov decision processes with well-defined semantics and decidable reachability and $\omega$-regular properties for important subclasses.
@@
Finally, formal modeling and verification of safety-critical systems has increasingly emphasized the integration of stochastic timing, uncertainty, and control [63, 10]. Tools like PRISM [35], MODEST [24], and Storm [15] support a subset of these features, often limited to MDP-based models. Controlled SANs provide a formalism that aligns more closely with the practical demands of such systems—offering compositionality, distributed semantics, and structured control—while retaining formal rigor and support for verification and synthesis. One of the most prominent tools for modeling and analyzing SAN-based systems is the Möbius Modeling Framework [14], which provides a compositional, extensible environment for performance and dependability analysis. Möbius has significantly influenced the modeling community and offers a natural foundation for extensions such as Controlled Stochastic Activity Networks (CSANs).
@@ 3 Nondeterministic Models
Controlled activity networks are a generalization of activity networks [48, 49], incorporating control mechanisms in a nondeterministic setting.
@@
Except for the inclusion of control actions, the structure of controlled activity networks is the same as the structure of activity networks [49]. As an example, consider the graphical representation of a controlled activity network as in Figure 1. $P1$, $P2$, $P3$, $P4$, $P5$, $P6$, $P7$, and $P8$ are places. $T1$, $T2$, $T3$, and $T4$ are timed activities, and $I1$, and $I2$ are instantaneous activities. $c1$, $c2$, $c3$, and $c4$ are control actions. $G1$ is an input gate with only one input, and $G2$ is an output gate with only one output. The enabling predicates, if any, and functions of these gates are indicated in a table called a “Gate Table” as depicted in Figure 1.
@@
Next, we define a notion of equivalence for controlled automata based on the concept of bisimulation, which is similar to the one proposed for automata [45, 3].
@@
The proof builds on the foundational result that extended Petri nets with inhibitor arcs can simulate nondeterministic Turing machines [23, 54, 65]. More specifically, the class of Generalized Stochastic Petri Nets (GSPNs) [40, 4] extends classical Petri nets with stochastic timed transitions and immediate transitions—closely resembling activity networks.
Controlled activity networks generalize GSPNs by integrating control actions and gate logic (e.g., standard, enabling, and inhibitor gates). These networks can include both instantaneous and stochastically timed activities, whose firing conditions may depend on the current marking and control policy.
By appropriately encoding the tape, state, and transition rules of a nondeterministic Turing machine into a controlled activity network with discrete places, inhibitor gates, and deterministic control actions, we can simulate any computable relation. Hence, the class of computable controlled automata coincides (up to isomorphism) with the class of automata realizable by such networks. For full details, see Appendix A. ∎
The above result shows that, in terms of behavioral equivalence, any computable controlled automaton can be represented using a controlled activity network comprising standard and inhibitor gates. However, such a representation may require substantial structural expansion—potentially leading to models that are less intuitive or harder to construct manually.
@@
Any computable controlled activity network is isomorphic to a network that uses only standard and inhibitor gates.
Remark 3.15.
While Generalized Stochastic Petri Nets (GSPNs) [40] offer a powerful formalism for performance modeling with stochastic timing, their expressiveness is inherently limited to bounded reachability and fixed transition structures. In contrast, Controlled SANs (CSANs) are Turing-complete and policy-driven, enabling the representation of arbitrary computable control behaviors and unbounded decision-making patterns. Thus, CSANs strictly subsume the modeling power of GSPNs, especially in contexts requiring adaptive control, dynamic policy synthesis, or integration with AI-driven strategies.
@@
We now define three types of policies for controlled automata by augmenting these models with finite-state automata, pushdown automata, and Turing machines [65] as follows.
@@
$L_{0}(S)\subseteq L_{F}(S)\subseteq L_{\text{Stack}}(S)\subseteq L_{\text{Tape}}(S)\subseteq L_{H}(S)$
@@
The next theorem shows that the expressive power of finite-memory control is inherently constrained: not every infinite regular family can be realized by a single controlled automaton.
Theorem 3.32.
@@
Definition 3.33.
@@
The next theorem shows that, at the regular level ($\Pi\in\{0,F\}$), this stronger form of uniform closure indeed holds via the standard disjoint-union and synchronous-product constructions.
Theorem 3.39.
@@
Determinization under Restricted Policy Types for Finite Words
We first record an analogous observation for finite words.
Theorem 3.44.
Let $S$ be a finite controlled automaton over finite words, and let the policy type be memoryless (type 0). Then there exists a deterministic controlled automaton $S_{\det}$ over the same input alphabet such that
$\{\,L_{\pi}(S):\pi\in\Pi_{0}(S)\,\}\;=\;\{\,L_{\pi^{\prime}}(S_{\det}):\pi^{\prime}\in\Pi_{0}(S_{\det})\,\}.$
In particular, for type 0 we may, without loss of closed-loop expressiveness over finite words, restrict attention to deterministic plants.
Proof.
For a fixed plant $S$, the set of memoryless policies $\Pi_{0}(S)$ is finite. Each policy $\pi\in\Pi_{0}(S)$ induces a closed-loop system “$S$ under $\pi$”, which is a finite-state automaton over the input alphabet, and hence the induced language $L_{\pi}(S)$ is regular. Therefore, for each $\pi\in\Pi_{0}(S)$ there exists a deterministic finite automaton $D_{\pi}$ recognizing $L_{\pi}(S)$.
We now combine these finitely many deterministic automata into a single deterministic plant $S_{\det}$ whose control alphabet contains one distinguished action for each policy $\pi\in\Pi_{0}(S)$, intuitively selecting the corresponding component $D_{\pi}$. A memoryless policy on $S_{\det}$ chooses one such action in the initial state and has no further essential influence on the run. By construction, for each $\pi\in\Pi_{0}(S)$ there is a corresponding memoryless policy $\pi^{\prime}\in\Pi_{0}(S_{\det})$ such that $L_{\pi}(S)=L_{\pi^{\prime}}(S_{\det})$, and conversely every closed-loop language of $S_{\det}$ under a memoryless policy arises in this way. This yields the claimed equality of families of languages. ∎
Remark 3.45.
For the finite-memory policy type (type $F$), the situation is more subtle. On a fixed plant $S$, the class of admissible finite-memory controllers may be infinite, so the above finitary construction (which explicitly enumerates all policies) does not directly apply. While each closed-loop language $L_{\pi}(S)$ is still regular and can be recognized by a deterministic finite automaton, it is not clear in general whether there always exists a single deterministic controlled automaton $S_{\det}$ whose finite-memory policies generate the same family of closed-loop languages as the original plant under all finite-memory policies. Clarifying this question for type $F$ appears to be an interesting open problem and is left for future investigation.
@@
Proof.
For $\Pi\in\{0,F\}$, controlled automata reduce to nondeterministic finite automata (NFAs). The emptiness problem for NFAs can be decided in polynomial time via a reachability check over the underlying state graph [29].
For $\Pi=\textit{Stack}$, the model corresponds to a pushdown automaton, and emptiness is known to be decidable using standard PDA algorithms [29].
For $\Pi\in\{\textit{Tape},H\}$, the policy mechanism can simulate a Turing machine, and the emptiness problem reduces to the halting problem, which is undecidable [65, 30]. ∎
@@
$L^{\omega}_{0}(S_{\omega})\subseteq L^{\omega}_{F}(S_{\omega})\subseteq L^{\omega}_{\text{Stack}}(S_{\omega})\subseteq L^{\omega}_{\text{Tape}}(S_{\omega})\subseteq L^{\omega}_{H}(S_{\omega})$
@@
We now lift our analysis to the setting of infinite words. Analogous to the finite-word hierarchy of language classes under various controller types, we obtain a similar strict hierarchy for $\omega$-languages under Büchi acceptance. The following theorem formalizes this structure.
@@
$\mathbb{L}^{\omega}_{0}\subset\mathbb{L}^{\omega}_{F}\subset\mathbb{L}^{\omega}_{\text{Stack}}\subset\mathbb{L}^{\omega}_{\text{Tape}}=\mathbb{L}^{\omega}_{H^{\text{comp}}}\subset\mathbb{L}^{\omega}_{H}\,.$
Proof.
The argument parallels that of Theorem 3.30 for finite-word languages, adapted to $\omega$-languages under Büchi acceptance using Lemma 3.55. Each inclusion corresponds to a strictly increasing level of controller power: memoryless $\subset$ finite-memory $\subset$ stack-augmented $\subset$ tape-augmented == computable history-dependent $\subset$ arbitrary history-dependent. Strictness follows from analogous examples where finite memory distinguishes finite phases, stacks enforce well-nested obligations, and tapes simulate arbitrary computations. Hence,
$\mathbb{L}^{\omega}_{0}\subset\mathbb{L}^{\omega}_{F}\subset\mathbb{L}^{\omega}_{\text{Stack}}\subset\mathbb{L}^{\omega}_{\text{Tape}}=\mathbb{L}^{\omega}_{H^{\text{comp}}}\subset\mathbb{L}^{\omega}_{H}\,.$
@@
As in the case of finite-word languages, the set of all $\omega$-regular languages realizable by finite-memory controllers on a fixed plant does not capture all infinite families of interest. In particular, even among countably infinite families of simple $\omega$-regular languages, realizability may require varying the underlying automaton. The following theorem makes this limitation precise.
Theorem 3.56.
@@
Proof.
This argument is the $\omega$-word analogue of Theorem 3.32. For each integer $k\geq 1$, define the $\omega$-regular language $W_{k}=(a^{k}b)^{\omega}$, i.e., the ultimately periodic infinite word consisting of blocks of $a^{k}$ followed by $b$, repeated forever. For distinct $k$, these $W_{k}$ are distinct $\omega$-regular languages.
@@
These results parallel Lemma 3.31 and Theorem 3.32 for finite words. The class $\mathbb{L}^{\omega}_{F}$ is strictly between $\mathbb{L}^{\omega}_{0}$ and $\mathbb{L}^{\omega}_{\mathrm{Stack}}$ and inherits the expressive power and closure properties of $\omega$-regular languages.
@@
Definition 3.58.
@@
Proof.
By Büchi’s theorem [8] and McNaughton’s determinization result [43], every $\omega$-regular language is accepted by a deterministic Muller automaton, and deterministic Muller automata are effectively closed under union, intersection, and complementation. Hence, $\mathrm{REG}^{\omega}$ enjoys full Boolean closure. ∎
@@
Proof.
Closure under union follows from the characterization of $\omega$-CFLs as finite unions of languages of the form $UV^{\omega}$ with $U,V$ context-free, as shown by Cohen and Gold [11]. Non-closure under intersection and complement is classical: there exist $\omega$-CFLs $L_{1},L_{2}$ such that $L_{1}\cap L_{2}$ is not $\omega$-context-free, implying, by De Morgan’s laws, failure of closure under complement (see also Staiger [66] and Finkel [21]). ∎
@@
Proof.
The class $\mathrm{RE}^{\omega}$ coincides with the family of analytic ($\Sigma^{1}_{1}$) subsets of $A^{\omega}$, known to be closed under finite union and intersection but not under complement (Kechris [33], Moschovakis [47]). ∎
@@
If, in addition, $\mathcal{S}_{\omega}$ is restricted to controlled deterministic parity automata (or any equivalent complement-closed $\omega$-regular formalism), then the same framework yields uniform closure under complementation as well, by replacing the parity condition with its dual on a fixed deterministic plant. Since we do not impose this structural restriction in general, Proposition 3.63 is stated only for union and intersection. It is well known that deterministic Büchi automata are not closed under complement, whereas deterministic parity or Rabin automata are. Hence, uniform complementation requires augmenting the acceptance structure beyond the Büchi property (e.g., to parity, Rabin/Street, or Muller conditions [60, 58, 22, 50]).